OpenStudy (anonymous):

If log sub b(3^b)=b/2,thenb=?

OpenStudy (anonymous):

\[If log sub b (3^b)=b/2,then b=?\]

jimthompson5910 (jim_thompson5910):

\[\large \log_{b}(3^b)=\frac{b}{2}\] \[\large \frac{\log_{10}(3^b)}{\log_{10}(b)}=\frac{b}{2}\] \[\large \frac{b*\log_{10}(3)}{\log_{10}(b)}=\frac{b}{2}\] \[\large \frac{\log_{10}(3)}{\log_{10}(b)}=\frac{1}{2}\] \[\large \frac{2*\log_{10}(3)}{\log_{10}(b)}=1\] \[\large \frac{\log_{10}(3^2)}{\log_{10}(b)}=1\] \[\large \frac{\log_{10}(9)}{\log_{10}(b)}=1\] \[\large \log_{b}(9)=1\] \[\large 9=b^1\] \[\large 9=b\] \[\large b=9\]